Organic Chemistry Question

Calculate the optical activity of a sample if it has 34% R, and 76% S-glutamic acid with a specific rotation of (-)8.9° for the R isomer.
Answer #1

enantiomelic Id-al encess, e.e. = 10+ T %100% = 134-761 41000% 1347761 = 4 20 x200 40 = 4268 30, (3) Hlutamic acid is present